OpenStudy (anonymous):

Find the two products below. Compare and contrast, in complete sentences, the similarities and differences of the two. (x + 6)(x − 6) and (x − 6)(x − 6)

OpenStudy (anonymous):

(x + 6)(x − 6) and (x − 6)(x − 6) --you have to times each figure with another figure. for example for the first two brackets, its x times x, which is x squared, the x times -6, then 6 times x and 6 times -6 . so the first product is (x + 6)(x − 6) =x^2-6x+6x+36 =x^2+36 ans... the other one is (x − 6)(x − 6) =x^2 -6x-6x+36 =x^2 -12x+36 ans....